### v2

@@ 1 Adversarial Attacks: Categories and Aims
2 Non-adversarial Shifts: From Synthetic to Real-World
Non-adversarial shifts, including natural data corruptions and perturbations, are rarely characterized. Only a few works in Computer Vision focus on such a type of model robustness assessment. They have introduced various benchmarking datasets carvalho2019machine to investigate the impact of naturally-shifted inputs on modern machine learning models. These benchmarks involve the introduction of corruptions and/or perturbations to standard (clean) datasets, allowing the assessment of different models’ resilience against corrupted images. For instance, ImageNet-C hendrycks2019benchmarking and ImageNet- hendrycks2019benchmarking serve as synthetic benchmarks, each focusing on distinct aspects of robustness: corruption and perturbation. These benchmarks decouple robustness benchmarking by applying image transformations to the original images from the ImageNet dataset. Corruptions involve significant changes in image statistics, offering a testing ground for out-of-distribution scenarios. The benchmark includes 15 types of corruption transformations selected from noise, blur, weather, and digital categories, with five severity levels controlling the degree of distribution shift. Perturbations, on the other hand, entail subtle transformations of original images, drawn from the same categories as corruptions but are designed to be more challenging to perceive visually. The perturbation benchmark aims to assess models’ performance in the face of subtle data distribution shifts.
Creating a comprehensive real-world robustness benchmark that incorporates systematic distribution shifts poses challenges compared to synthetic benchmarks. The complexity arises from the multifaceted variations that can simultaneously occur in real-world data. For instance, images featuring the same object may be captured with different cameras, from diverse viewpoints, in various locations, and under different weather conditions. However, the issue raised in taori2020measuring underscores the necessity of real-world benchmarking datasets. ImageNetV2 recht2019imagenet , generated by replicating the original ImageNet data collection process, serves as an example. While explicit definitions of distribution shifts may be elusive, classifiers trained on ImageNet and assessed on ImageNetV2 demonstrate reduced performance, indicating inherent natural distribution shifts in the ImageNetV2 dataset. A parallel observation holds for ImageNet-Renditions (ImageNet-R) hendrycks2021many , encompassing renditions such as paintings, sculptures, and embroidery for ImageNet classes. This dataset introduces images with distinctly different textures and local image statistics, inducing a distribution shift relative to ImageNet. The StreetView StoreFronts (SVSF) hendrycks2021many dataset contains business storefront images. In the dataset, there is a wide variation in location, date of creation, and camera properties which cause sensor-induced distribution shifts. The real-world benchmark DeepFashion Remixed (DFR) hendrycks2021many leverages descriptive information to make systematic changes in object occlusion, zoom, orientation, and scale. Temporal Perturbations shankar2021image are natural perturbations that are deduced from small changes occurring in assembled sets of contiguous video frames that appear perceptually similar to humans, but might produce inconsistent predictions for ML models. All these perturbations result in a condition where the distribution of the test set differs from the one of the training set.
Unlike adversarial attacks that aim to exploit model vulnerabilities, non-adversarial or natural perturbations serve as synthetic or newly created testing datasets for evaluating models in an out-of-distribution manner, simulating shifts from their original training distribution. These perturbations effectively highlight the models’ fragility to naturally occurring data distortions. However, the absence of performance degradation on these datasets does not necessarily imply robustness, as they may suffer from selection and sample bias. For instance, when creating datasets with data augmentation techniques like changing image brightness, the use of brightness-based corruptions might not be indicative of model robustness, as the training data distribution already encompasses various brightness levels. In this case, non-adversarial robustness represents how well the model maintains its predictive performance under foreseeable input shifts (i.e, variation of brightness level).
@@ 3 DL Software Testing: Test-Driven Model Verification
Deep Learning (DL) software testing has emerged as a class of model verification techniques which bridges the gap between adversarial attacks and non-adversarial data shifts. As such, it differs from adversarial attacks in the sense that the methods are not used to improve a model resilience, but rather to evaluate a certain number of properties (associated with robustness) after the model was refined and/or trained. Hence, it does not act on the model but rather aims at verifying it. DL software testing renovates the conventional software testing methods, including test oracle identification, test adequacy evaluation, and test input generation, to be specialized for finding unrestricted adversarial examples that expose the target DNN brittleness. The goal is to construct systematic approaches that produce synthetic inputs, representing both major and minor naturally-occurring conditions to reveal potential incorrect behaviors. This involves optimizing the search over the space of transformed data as opposed to randomly sampling non-adversarially distorted inputs (i.e., non-adversarial shifts). Therefore, the designed test cases should cover effectively a broader set of input-output mappings than iid evaluations to complement the conventional statistical testing.
@@ 4 Empirical Methods and Their Limitations
Empirical robustness assessment provides a practical, but valuable approach to complement the iid performance by statistical scoring of the model robustness in the face of unexpected data changes. On one extreme, adversarial examples provide worst-case perturbations with a defined expectation that models should remain performant under small, bounded perturbations. On the other extreme, naturally-occurring distortions that can produce corrupted entries under various sources of noise require a controlled severity procedure to avoid generating unrecognizable inputs, even to humans. Robustness assessment needs to better account for expected model behavior between these extremes, emphasizing the necessity to identify data changes’ preconditions under which a model’s performance is not expected to deteriorate. Especially for robustness to natural and non-adversarial perturbations, characterizing the type of data the model might encounter is crucial. This cannot be done algorithmically based on the analysis of the model’s behaviors or in-data distribution features. It is vital to find better interfaces for domain knowledge in modern ML pipelines. Domain-aware testing methods are needed to thoroughly test models on application-specific tasks and check that performance on these tasks remains stable beyond the collected data samples used as in-distribution data. Designing stress tests underspecificationMLmodel well-matched to applied requirements and providing good “coverage” of potential failure modes is a significant challenge that requires incorporating domain knowledge. For intance, braiek2022physics proposed to leverage physics first principles and system-related design properties in the form of input-output sensitivities to generate invariance or directional expectations tests, i.e., the prediction should remain the same or should change direction under the introduction of input changes, respectively. Thus, the failed tests expose the inconsistencies of trained aircraft performance models with the domain-specific knowledge.
@@ 1 Data-Centric Amelioration Strategies
Data augmentation aims to enhance the diversity among the training samples and enlarge the size of distinctive instances through the ERM. It involves data transformations that derive new synthetic inputs supporting the model generalization by promoting the learning of feature representations insensitive to irrelevant noises or naturally-occuring distortions. For instance, basic image augmentations can be pixel-value transformations like brightness and contrast modifiers, or geometric transformations such as random rotation, translation, and mirroring. The severity of such distortions should be tuned to prevent significantly altering the data distribution but ensures a beneficial regularization effect on the model. Other augmentation techniques cutout ; lee2020smoothmix systematically eliminate information from the training inputs to increase the robustness of the model against occlusions and information loss in general. The cutout augmentation cutout works similarly to dropout, by nullifying activations of input neurons associated with arbitrary input image patches. As a result, the network learns not to heavily rely on any specific feature of the data. In order to step beyond the null patches, Cutmix yun2019cutmix propose to instead inject patches that have been cut out from other images in the training dataset. In a higher level of sophistication, we can find the mixup augmentation zhang2017mixup that relies on linear mixing of training images and labels to encourage more stable predictions on data outside the training distribution. Mixup-derived cutout methods have contributed to continued enhancements in corruption robustness. PuzzleMix kim2020puzzle improves upon Mixup by optimizing the amount of information retained in the original pairs of images while preserving saliency information during the mixing procedure. To alleviate sharp edge effects, Smoothmix lee2020smoothmix creates a smooth blending mask by randomly sampling mask shape and associated shape parameters (e.g., a square mask with sampled dimensions), then, the pairs of images are combined using the relative proportions of each image derived from the mask.
@@
In order to create a complex augmentation, multiple simple augmentations can be chained together costante2015exploring . In a more recent method, AugMix hendrycks2019augmix further improved corruption robustness of models by using a weighted combination of augmentations while enforcing consistent embeddings of the augmented images. Even though studies cutout ; dodge2017can ; lee2020smoothmix emphasize the importance of combining multiple augmentations to achieve good generalization, it is a non-trivial task to find the optimal combination. In some cases, inappropriate combinations can cause a degradation in model generalization such as in  dodge2017can for DeepFashion Remixed benchmark dataset. Most software model testing approaches deepxplore ; deeptest ; feng2020deepgini implement test input prioritization and systematic test generation that uses data transformations similar to augmentation techniques, but searches for more diverse transformed data with fault-revealing abilities. The generated synthetic datasets, especially those resulting in false predictions, can be added to augment the training data. Indeed, their associated labels can be deduced based on the predefined metamorphic relationship of the input-output transformation deeptest ; feng2020deepgini , or based on the majority voting deepxplore to automatically generate labels for the generated test inputs. Thus, no manual labeling is required to perform a conventional supervised learning of the model using the original dataset plus a selection of synthetically-produced inputs. This retraining often results in fixing the majority of the revealed erroneous behaviors while preserving or hopefully improving its performance.
@@
However, data augmentation has been criticized because it heavily relies on rule-based data transformations, which can lead to semantically non-meaningful combinations that are not compatible with real-world data taori2020measuring , e.g., a horse with a patch covering the head of a frog]. Therefore, generative-based augmentations are an important initiative to produce uniquely synthetic images using generative models trained on the in-data distribution, but they can override the inherent characteristics of samples in order to cover underrepresented input regions. For instance, a style transfer variant of generative adversarial networks (GAN) GAN proposed in taori2020measuring has been shown to facilitate robustness against out-of-distribution when used as an augmentation technique hendrycks2021many . GAN-based solutions are also proven useful in enhancing the model’s immunity to adversarial attacks abdelaty2021gadot ; wang2019direct ; sun2019enhancing . In particular, they are employed to generate adversarial samples abdelaty2021gadot , perturbations wang2019direct , and boundary samples sun2019enhancing to defend the networks against adversarial attacks.
@@ 2 Model-Centric Optimization Methods
Transfer learning is a machine learning technique that involves transferring knowledge learned by the model in one task to improve another, related task. Transfer learning is based on the notion that DNN learns feature representations gradually from simple, task-agnostic features (e.g., lines) to complex, task-specific features (e.g., nose or ears for face recognition task) that can be fully or partially transferred to another problem. A common way to implement transfer learning is to pretrain a DNN on a large and diverse dataset and then use the first n pretrained layers (i.e., known as feature extractor) as an initialization for a new DNN that is then trained on a new dataset. Transfer learning methods are many, and the selected method determines the efficiency of the knowledge transfer from the pretrained model to the new one. In regards to robustness, it has been demonstrated that in certain settings it can increase the models OOD robustness, both in adversarial and non-adversarial settings hendrycks2019using . Intuitively, the use of a small supervised dataset with an overparameterized neural network may provide high IID performance, but there is no guarantee regarding its behavior against out of distribution inputs due to shortcut learning and under-fitted neurons. This risk of shortcut learning and under-fitted neurons can be diminished by using pretrained models. Furthermore, the advances in self-supervised learning (SSL) for representation learning enables the exploitation of massive unlabeled target datasets to train models that solve an auxiliary, pretext task to learn rich feature representations, e.g., learning to distinguish between images chen2020simple ; caron2021emerging . Transfer learning with SSL sun2019unsupervised ; sun2020test involves using these self-learned models as pretrained feature extractors.
@@
Kurakin et al. IFGSM demonstrates that adversarial training can be applied in massive datasets like ImageNet, showing an increase in robustness for one-step adversarial methods. As a result, several adversarial training approaches have been proposed to maximize the robustness of a model during deployment by adjusting the decision boundaries in response to adversarial perturbations computed based on training data.
Adversarial training with one-step attacks is prone to overfitting, especially with large datasets. The overfitting problem hinders the convergence to a global minimum, which makes the adversarially-trained models vulnerable to black-box attacks. Tramer et al. tramer2017ensemble leverage multiple pre-trained models to produce adversarial examples for AT. They aim to remove the tight coupling between the on-training model and the adversarial attack, which could discredit the adversarial examples in the long run. Mummadi et al. mummadi2019defending propose to compute the gradients w.r.t a batch of training inputs rather than one single input, then, the gradients are merged to create a shared perturbation for the entire batch of inputs. Once these shared perturbations have been iteratively refined and clipped, they are used to augment training data during AT.
@@ 3 Post-training Model Enhancements
The strategies, which involve changing training data or updating the model, require retraining or fine-tuning, which is costly since modern neural networks are complex. Moreover, pretrained models can sometimes be obtained from a third party or training data may be confidential, making access to the whole in-data distribution challenging. Importantly, there is a disconnection between the assessment techniques and the improvement processes for model robustness. There is no guarantee that the new optimized or upgraded model will respond correctly to the counterexamples discovered during the assessment using most amelioration techniques. Robust models are often evaluated on adversarial examples or shifted inputs generated by assessment methods. It is important to show that the amelioration strategy does not overfit on counterexamples, but it is often overlooked how successful it is at fixing the revealed erroneous behaviors. Therefore, researchers have proposed post-training, model-level repair of DNNs, i.e., repair through the modification of the weights of an already trained model. In conventional software systems, these narrowly-defined modifications are called patches. When it comes to deep learning, the patch aims to fix the model’s erroneous behaviors against a specific subset of counterexamples. Although the high complexity of modern ML models makes it difficult to perform provable model repairs (i.e., the counterexample provided will always be corrected), heuristic repair strategies that do not guarantee to repair all observed counterexamples still have benefits when compared to retraining or fine-tuning the model. Indeed, the original neural network likely went through training and functions properly on most inputs, so applying small, less impactful but precise changes to fix brittleness makes sense. CAusality-based REpair (CARE) sun2022causality identifies the “‘guilty” neurons (i.e., the ones that caused the false prediction) using causality-based fault localization, then, modifies the weights of these identified neurons to reduce the misbehaviors. Arachne sohn2023arachne is similar to CARE, but ensures no disturbance of the model’s correct behaviors. Indeed, it uses differential evolution to generate effective patches for the localised weights that can fix specific mispredictions of a DNN without drastically reducing its clean accuracy. GenMuNN wu2022genmunn ranks the weights based on their influences on the model’s predictions. It then generates mutants using the computed ranks and evolves them using genetic algorithms to increase the chances of finding mutants that satisfy the stopping criteria. By tracking the training history, NeuRecover tokui2022neurecover finds weights that have changed significantly over time. A weight becomes a subject for repair when it is no longer contributing to correct predictions in the earlier stage of training but is leading to incorrectly-predicted inputs. Similarly to NeuRecover, I-Repair henriksen2022repairing modifies localised weights to influence predictions for a specific set of fault-revealing inputs while minimising the impact on correct predictions. NNrepair NNRepair uses fault localisation to identify suspicious weights, and then uses constraint solving to modify them marginally. There are provable model repairs like PRDNN sotoudeh2021provable , REASSURE fu2021sound , and Minimal Modifications of DNNs (MMDNN) goldberger2020minimal . However, these approaches are not scalable to large DNNs, only support ReLU activation, and often do not support polytope repair or multilayer repairs.

### v3

@@ 3 DL Software Testing: Test-Driven Model Verification
Deep Learning (DL) software testing has emerged as a class of model verification techniques which bridges the gap between adversarial attacks and non-adversarial data shifts. As such, it differs from adversarial attacks in the sense that the methods are not used to improve a model resilience, but rather to evaluate a certain number of properties (associated with robustness) after the model was refined and/or trained. Hence, it does not act on the model but rather aims at verifying it. DL software testing renovates the conventional software testing methods, including test oracle identification, test adequacy evaluation, and test input generation, to be specialized for finding unrestricted adversarial examples that expose the target DNN brittleness. The goal is to construct systematic approaches that produce synthetic inputs, representing both major and minor naturally-occurring conditions to reveal potential incorrect behaviors. This involves optimizing the search over the space of transformed data as opposed to randomly sampling non-adversarially distorted inputs (i.e., non-adversarial shifts). Therefore, the designed test cases should cover effectively a broader set of input-output mappings than i.i.d. evaluations to complement the conventional statistical testing.
@@ 4 Empirical Methods and Their Limitations
Empirical robustness assessment provides a practical, but valuable approach to complement the i.i.d. performance by statistical scoring of the model robustness in the face of unexpected data changes. On one extreme, adversarial examples provide worst-case perturbations with a defined expectation that models should remain performant under small, bounded perturbations. On the other extreme, naturally-occurring distortions that can produce corrupted entries under various sources of noise require a controlled severity procedure to avoid generating unrecognizable inputs, even to humans. Robustness assessment needs to better account for expected model behavior between these extremes, emphasizing the necessity to identify data changes’ preconditions under which a model’s performance is not expected to deteriorate. Especially for robustness to natural and non-adversarial perturbations, characterizing the type of data the model might encounter is crucial. This cannot be done algorithmically based on the analysis of the model’s behaviors or in-data distribution features. It is vital to find better interfaces for domain knowledge in modern ML pipelines. Domain-aware testing methods are needed to thoroughly test models on application-specific tasks and check that performance on these tasks remains stable beyond the collected data samples used as in-distribution data. Designing stress tests underspecificationMLmodel well-matched to applied requirements and providing good “coverage” of potential failure modes is a significant challenge that requires incorporating domain knowledge. For intance, braiek2022physics proposed to leverage physics first principles and system-related design properties in the form of input-output sensitivities to generate invariance or directional expectations tests, i.e., the prediction should remain the same or should change direction under the introduction of input changes, respectively. Thus, the failed tests expose the inconsistencies of trained aircraft performance models with the domain-specific knowledge.
@@ 1 Data-Centric Amelioration Strategies
Data augmentation aims to enhance the diversity among the training samples and enlarge the size of distinctive instances through the ERM. It involves data transformations that derive new synthetic inputs supporting the model generalization by promoting the learning of feature representations insensitive to irrelevant noises or naturally-occuring distortions. For instance, basic image augmentations can be pixel-value transformations like brightness and contrast modifiers, or geometric transformations such as random rotation, translation, and mirroring. The severity of such distortions should be tuned to prevent significantly altering the data distribution but ensures a beneficial regularization effect on the model. Other augmentation techniques cutout ; lee2020smoothmix systematically eliminate information from the training inputs to increase the robustness of the model against occlusions and information loss, in general. The cutout augmentation cutout works similarly to dropout, by nullifying activations of input neurons associated with arbitrary input image patches. As a result, the network learns not to heavily rely on any specific feature of the data. In order to step beyond the null patches, Cutmix yun2019cutmix propose to instead inject patches that have been cut out from other images in the training dataset. In a higher level of sophistication, we can find the mixup augmentation zhang2017mixup that relies on linear mixing of training images and labels to encourage more stable predictions on data outside the training distribution. Mixup-derived cutout methods have contributed to continued enhancements in corruption robustness. PuzzleMix kim2020puzzle improves upon Mixup by optimizing the amount of information retained in the original pairs of images while preserving saliency information during the mixing procedure. To alleviate sharp edge effects, Smoothmix lee2020smoothmix creates a smooth blending mask by randomly sampling mask shape and associated shape parameters (e.g., a square mask with sampled dimensions), then, the pairs of images are combined using the relative proportions of each image derived from the mask.
@@
To create a complex augmentation, multiple simple augmentations can be chained together costante2015exploring . AugMix hendrycks2019augmix improves corruption robustness of models by using a weighted combination of augmentations while enforcing consistent embeddings of the augmented images. Even though studies cutout ; dodge2017can ; lee2020smoothmix emphasize the importance of combining multiple augmentations to achieve good generalization, it is a non-trivial task to find the optimal combination. In some cases, inappropriate combinations can cause a degradation in model generalization such as in  dodge2017can for the DeepFashion Remixed benchmark dataset. Most software model testing approaches deepxplore ; deeptest ; feng2020deepgini implement test input prioritization and systematic test generation that uses data transformations similar to augmentation techniques, but searches for more diverse transformed data with fault-revealing abilities. The generated synthetic datasets, especially those resulting in false predictions, can be added to augment the training data. Indeed, their associated labels can be deduced based on the predefined metamorphic relationship of the input-output transformation deeptest ; feng2020deepgini , or based on the majority voting deepxplore to automatically generate labels for the generated test inputs. Thus, no manual labeling is required to perform a conventional supervised learning of the model using the original dataset plus a selection of synthetically-produced inputs. This retraining often results in fixing the majority of the revealed erroneous behaviors while preserving or hopefully improving its performance.
@@
However, data augmentation has been criticized because it heavily relies on rule-based data transformations, which can lead to semantically non-meaningful combinations that are not compatible with real-world data taori2020measuring , e.g., a horse with a patch covering the head of a frog. Therefore, generative-based augmentations are an important initiative to produce uniquely synthetic images using generative models trained on the in-data distribution, but they can override the inherent characteristics of samples in order to cover underrepresented input regions. For instance, a style transfer variant of generative adversarial networks (GAN) GAN proposed in taori2020measuring has been shown to facilitate robustness against out-of-distribution when used as an augmentation technique hendrycks2021many . GAN-based solutions are also proven useful in enhancing the model’s immunity to adversarial attacks abdelaty2021gadot ; wang2019direct ; sun2019enhancing . In particular, they are employed to generate adversarial samples abdelaty2021gadot , perturbations wang2019direct , and boundary samples sun2019enhancing to defend the networks against adversarial attacks.
@@ 2 Model-Centric Optimization Methods
Transfer learning is a ML technique that involves transferring knowledge learned by the model in one task to improve another, related task. Transfer learning is based on the notion that DNN learns feature representations gradually from simple, task-agnostic features (e.g., lines) to complex, task-specific features (e.g., nose or ears for face recognition task) that can be fully or partially transferred to another problem. A common way to implement transfer learning is to pretrain a DNN on a large and diverse dataset and then use the first n pretrained layers (i.e., known as feature extractor) as an initialization for a new DNN that is then trained on a new dataset. Transfer learning methods are many, and the selected method determines the efficiency of the knowledge transfer from the pretrained model to the new one. In regards to robustness, it has been demonstrated that in certain settings it can increase the models OOD robustness, both in adversarial and non-adversarial settings hendrycks2019using . Intuitively, the use of a small supervised dataset with an overparameterized neural network may provide high ID performance, but there is no guarantee regarding its behavior against out of distribution inputs due to shortcut learning and under-fitted neurons. This risk of shortcut learning and under-fitted neurons can be diminished by using pretrained models. Furthermore, the advances in self-supervised learning (SSL) for representation learning enables the exploitation of massive unlabeled target datasets to train models that solve an auxiliary, pretext task to learn rich feature representations, e.g., learning to distinguish between images chen2020simple ; caron2021emerging . Transfer learning with SSL sun2019unsupervised ; sun2020test involves using these self-learned models as pretrained feature extractors.
@@
Kurakin et al. IFGSM demonstrates that AT can be applied in massive datasets like ImageNet, showing an increase in robustness for one-step adversarial methods. As a result, several adversarial training approaches have been proposed to maximize the robustness of a model during deployment by adjusting the decision boundaries in response to adversarial perturbations computed based on training data.
AT with one-step attacks is prone to overfitting, especially with large datasets. The overfitting problem hinders the convergence to a global minimum, which makes the adversarially-trained models vulnerable to black-box attacks. Tramer et al. tramer2017ensemble leverage multiple pre-trained models to produce adversarial examples for AT. They aim to remove the tight coupling between the on-training model and the adversarial attack, which could discredit the adversarial examples in the long run. Mummadi et al. mummadi2019defending propose to compute the gradients w.r.t a batch of training inputs rather than one single input, then, the gradients are merged to create a shared perturbation for the entire batch of inputs. Once these shared perturbations have been iteratively refined and clipped, they are used to augment training data during AT.
@@ 3 Post-training Model Enhancements
The strategies, which involve changing training data or updating the model, require retraining or fine-tuning, which is costly since modern neural networks are complex. Moreover, pretrained models can sometimes be obtained from a third party or training data may be confidential, making access to the whole in-data distribution challenging. Importantly, there is a disconnection between the assessment techniques and the improvement processes for model robustness. There is no guarantee that the new optimized or upgraded model will respond correctly to the counterexamples discovered during the assessment using most amelioration techniques. Robust models are often evaluated on adversarial examples or shifted inputs generated by assessment methods. It is important to show that the amelioration strategy does not overfit on counterexamples, but it is often overlooked how successful it is at fixing the revealed erroneous behaviors.
@@
Therefore, researchers have proposed post-training, model-level repair of DNNs, i.e., repair through the modification of the weights of an already trained model. In conventional software systems, these narrowly-defined modifications are called patches. When it comes to deep learning, the patch aims to fix the model’s erroneous behaviors against a specific subset of counterexamples. Although the high complexity of modern ML models makes it difficult to perform provable model repairs (i.e., the counterexample provided will always be corrected), heuristic repair strategies that do not guarantee to repair all observed counterexamples still have benefits when compared to retraining or fine-tuning the model. Indeed, the original neural network likely went through training and functions properly on most inputs, so applying small, less impactful but precise changes to fix brittleness makes sense. CAusality-based REpair (CARE) sun2022causality identifies the “‘guilty” neurons (i.e., the ones that caused the false prediction) using causality-based fault localization, then, modifies the weights of these identified neurons to reduce the misbehaviors. Arachne sohn2023arachne is similar to CARE, but ensures no disturbance of the model’s correct behaviors. Indeed, it uses differential evolution to generate effective patches for the localised weights that can fix specific mispredictions of a DNN without drastically reducing its clean accuracy. GenMuNN wu2022genmunn ranks the weights based on their influences on the model’s predictions. It then generates mutants using the computed ranks and evolves them using genetic algorithms to increase the chances of finding mutants that satisfy the stopping criteria. By tracking the training history, NeuRecover tokui2022neurecover finds weights that have changed significantly over time. A weight becomes a subject for repair when it is no longer contributing to correct predictions in the earlier stage of training but is leading to incorrectly-predicted inputs. Similarly to NeuRecover, I-Repair henriksen2022repairing modifies localised weights to influence predictions for a specific set of fault-revealing inputs while minimising the impact on correct predictions. NNrepair NNRepair uses fault localisation to identify suspicious weights, and then uses constraint solving to modify them marginally. There are provable model repairs like PRDNN sotoudeh2021provable , REASSURE fu2021sound , and Minimal Modifications of DNNs (MMDNN) goldberger2020minimal . However, these approaches are not scalable to large DNNs, only support ReLU activation, and often do not support polytope repair or multilayer repairs.